# Audio Personas: Augmenting Social Perception via Body-Anchored Audio Cues

### v1

@@ Abstract.
We introduce Audio Personas, enabling users to ”decorate” themselves with body-anchored sounds in audio augmented reality. Like outfits, makeup, and fragrances, audio personas offer an alternative yet dynamic channel to augment face-to-face interactions. For instance, one can set their audio persona as rain sounds to reflect a bad mood, bee sounds to establish personal boundaries, or a playful ”woosh” sound to mimic passing by someone like a breeze. To instantiate the concept, we implemented a headphone-based prototype with multi-user tracking and audio streaming. Our formative study with designers revealed that audio personas were preferred in public and semi-public-private spaces for managing social impressions (e.g., personality) and signaling current states (e.g., emotions). Our preregistered in-lab study with 64 participants showed that audio personas influenced how participants formed impressions. Individuals with positive audio personas were rated as more socially attractive, more likable, and less threatening than those with negative audio personas.
@@ 1. Introduction
While people can easily modify their social images online, face-to-face interactions lack such flexibility. Physical adornments like makeup and clothing remain relatively static, limited in their ability convey changing social messages. As such, there is a growing trend toward integrating digital elements, traditionally limited to virtual interactions, into real-world settings. One prominent example is the use of visual filters, popularized through social media, are now adopted in mobile augmented reality (mobile AR). By anchoring visual effects such as face masks or virtual jewelries to one’s body, it alters how one is socially perceived by others (Häkkilä et al., 2017; Roinesalo et al., 2016).
@@
We introduce the concept of audio personas111Here, a persona refers to the aspect of someone’s character that is presented to or perceived by others (Oxford English Dictionary, 2023; Goffman, 2023)., body-anchored audio cues to augment social perception in face-to-face interactions. Sound effects are commonly used in online games to provide immediate feedback on user actions and enhance engagement (Liljedahl, 2011; Lipscombc, 2013; Kenwright, 2020). However, it’s potential use face-to-face social contexts is largely unexplored. In fact, psychology research has shown that mere exposure to emotion-evoking sounds can influence how people perceive the social attractiveness and emotional states of others in photographs (May and Hamilton, 1980; Hanser et al., 2015; Logeswaran and Bhattacharya, 2009). With the rise of spatial audio, the ability to spatially anchor sounds to individuals—similar to how visual filters personalize appearance—creates a novel means of personalized expression through sounds.
@@
Figure 1 visualizes a few examples of audio personas: one can choose to surround themselves with bee sounds to set personal boundaries, a ”woosh” sound to mimic passing by like a breeze for fun and playfulness, or keyboard-clicking sounds to signal that they are in a busy state. To instantiate this concept, we first built a headphone-based prototype that utilizes multi-person tracking and multi-channel audio streaming. The user anchors sounds to their body and projects them to others in audio augmented reality.
We conducted two user studies to explore the concept of audio personas from two perspectives: (1) when users design and project audio personas, i.e., senders, and (2) when users listen to the audio personas of others, i.e., receivers.
Study #1 involved 8 designers to gain an initial understanding of the social intentions and contexts for using audio personas. We identified intended use cases for audio personas in public and semi-public-private spaces as a way to manage social impressions (e.g., personality) and signal current states (e.g., emotions). We also elicited the creations of 48 audio persona designs and found that audio personas being body-anchored is a key design feature. Additionally, we identified diverse design patterns in localization, triggers, dynamics and content of audio personas.
@@
Study #2 built upon the identified use cases, examining how audio personas influence social impression and perceived emotional states from the receiver’s perspective. In this preregistered in-lab study (n = 64), we focused on key design factors of audio personas (i.e., body-anchored, audio valence). Participants interacted with two research assistants, while hearing the research assistants’ audio personas or hearing sounds anchored to objects in the space. We found that participants were more likely to incorporate audio personas, body-anchored audio cues, into their impressions of others compared to the object-anchored audio. Moreover, participants perceived the person with a positive audio persona as more socially attractive, likable and less threatening than the person with a negative audio persona. These findings align with the theoretical grounding that social perception is a multisensory process (Livesley and Bromley, 1973; Workman and Johnson, 1991). The spatial alignment of audio with a person, as introduced by audio personas, enhances the likelihood of the sound being perceived as relevant cues of that individual in social interactions.
@@
Through a formative study with designers, we explored the social intentions and contexts for audio persona usage. This study also uncovered design possibilities for audio personas and identified key design patterns.
@@ 2.3. Dynamic Auditory Cues to Influence Social Perception
Our bodies generate sounds both within themselves (e.g., breathing, heartbeat) and through interactions with the environment (e.g., tapping sounds). Tajadura-Jiménez et al. introduced the concept of ”sonic self-avatar,” emphasizing that self-representations in mixed reality environments should incorporate body-centered sounds (Tajadura-Jiménez et al., 2018; Tajadura and Västfjäll, 2008). When these body-centered sounds are altered, they can influence users’ mental representations of their bodies (Tajadura-Jiménez et al., 2015b; D’Adamo et al., 2024; Tonetto et al., 2014), perception of objects (Furfaro et al., 2015), and even gaits and motion (Tajadura-Jiménez et al., 2015a). For example, amplifying the high frequencies of footstep sound in walking can create the self-perception of having a thinner body (Tajadura-Jiménez et al., 2015a). These audio-based sensory feedback devices were further explored in the context of physical activity and rehabilitation (Ley-Flores et al., 2024; Turmo Vidal et al., 2024; Gomez-Andres et al., 2020).
@@ 3.2. Audio Rendering and Multi-channel Audio Streaming
4. Study #1: User Perception and Designs of Audio Personas
What do people want to express with audio personas, and how do they use audio persona to convey information? To explore user perception and the design possibilities of audio personas, we conducted a formative study.
We chose to study participants with prior professional experience in audio. This sample selection allows us to understand the potential contexts and intentions behind audio persona usage from a user perspective, as well as elicit designs with expert insights.
4.1. Methods
4.1.1. Participants
4.1.2. Procedure
Participants were first introduced to the concept of audio personas. They put on the prototype system, as described in Sec. 3, and interacted with an experimenter to experience three examples of audio personas: (1) a bee sound when approaching the experimenter, (2) a ”woosh” sound when the experimenter walked past, and (3) exaggerated stomping sounds from the experimenter. These demos intended to provide inspiration, and participants were encouraged to think beyond these audio persona designs.
The study took approximately 75 - 90 minutes, and participants were compensated $35 for their time. The study was approved by the Institutional Review Board (IRB # 71992).
4.1.3. Analysis
We video-recorded the sessions, transcribed the data, and organized it based on the guiding interview questions, as described in the previous section and shown in Figure 3. Two experimenters participated in coding the transcriptions and designs, and we identified common themes through inductive thematic analysis (Braun and Clarke, 2006). The themes identified through thematic analysis are reflected in the headings of the following subsections.
4.2. Finding #1: Social Intentions of Using Audio Personas
4.2.1. Impression management
In the designs created by participants, 21 out of 48 audio personas intended to manage impressions, focusing on influencing how others perceive users. A wide range of audio content with varying valence was used and Figure 5 shows a selection of them. For instance, P7 designed an audio persona that features auditory accessories with a jingling bell sound to reflect personal jewelry taste and personality, as shown in (Figure 5a). P5 chose to augment their footstep sounds with a trombone sound, showing the love of SpongeBob (Figure 5b). P4 designed an audio persona that emits a dog barking sound whenever they carry their dog in a pouch, signaling that they are a pet owner and reminding others of the dog’s presence (Figure 5e).
4.2.2. Signaling current states
4.2.3. Limitations in delivering complex information
4.3. Finding #2: Social Context for Using Audio Personas
4.3.1. Intended for public and semi-public-private spaces
4.3.2. Less useful in private space
4.4. Finding #3: Design Patterns of Audio Personas
4.4.1. Localizations
We traced the sketches of audio persona locations marked by participants on the design worksheets and overlaid them to identify design patterns, as shown in Figure 8. One key observation across all 48 designs was that designers consistently anchored the audio persona to the user’s body, aligning the sound’s spatial location with the person’s. There was variability in the specific body parts where the audio personas were attached. 25 out of 48 designs were anchored to the upper body of the users, such as to the head (n = 11), hands (n = 4), and back (n = 3). 15 designs were placed around the lower body of the users, predominantly on the feet (n = 11). In 12 of 48 designs, participants circled the whole body instead of specifying a center origin, indicating they intended the audio personas to be heard around the body without a precise origin. 8 designs placed audio personas on multiple body parts.
4.4.2. Triggers
4.4.3. Dynamics
4.5. Summary of Findings in Study #1
In Study #1, we explored social context and intended social messages for using audio personas by working with designer participants. We identified potential use cases in managing social impressions (e.g., personality & traits, identity & belonging) and signaling current states (e.g., emotions) but are limited in communicating complex information (e.g., ideas & opinions). Participants preferred using audio personas in public (e.g., street, transport) and semi-private settings (e.g., workspace, educational institutions) to form bonds with people they do not know well, but found it less useful in private spaces. Analysis of 48 participant-generated audio personas revealed that anchoring sounds to the body is a key design feature. Additionally, a wide range of audio content with varied valence was used, and there are diverse design possibilities for localization, triggers, and dynamics.
5. Study #2: How Audio Persona Impacts Social Perception
Study #1 uncovered the intended use of audio personas to manage social impressions and signal current states. To explore whether audio personas are effective in influencing these dimensions of social perception, we need to examine from the receiver’s perspective. Does listening to others’ audio personas change how one perceive others?
In Study #2, 64 participants were invited to the lab for an icebreaker with two research assistants, with or without hearing the research assistants’ audio personas. As Study #1 identified body-anchored as a key design feature for audio persona and found that audio valence conveys distinct messages, Study #2 was designed around these two factors to examine the effect of audio personas. We hypothesized that hearing sounds anchored to a person (i.e., audio persona) would facilitate impression formation with the considerations of sounds. Furthermore, we hypothesized that the valence of the sounds would shape the direction of impressions and the perceived emotional states of others.
@@ 5.1. Condition Design
Audio personas being body-anchored was a key design feature observed in Study #1. Therefore, to understand the downstream impact of audio personas, it’s essential to understand how this spatial anchoring plays a role.
Specifically, participants were divided into two groups. Half of the participants were assigned to the audio persona group (Figure 10a, 10b), in which participants heard sounds anchored to each research assistants. As the research assistant walked past participants (e.g., moving from left to right or vice versa), their audio persona shifted in locations corresponding to their movements.
In the control group, participants interacted with research assistants without hearing their audio personas. Instead, participants heard sounds anchored to objects (i.e., speakers) in the space. We chose to include sounds in the control group as previous research showed that the mere presence of audio can already influence social emotional perception and behavior (Tajadura-Jiménez et al., 2011; Logeswaran and Bhattacharya, 2009; Carroll and Young, 2005). This control design thus better examine the exact effect of ”body-anchored” audio. Anchoring sound to objects is also commonly used in audio augmented reality (McGookin et al., 2009; Yang et al., 2019; Sikora et al., 2018) and the auditory experience in this group is similar to hearing sounds played from a physical speaker. The research assistants’ movements would not affect the sound.
@@
In Study #1, participants leveraged sounds with varying valences in audio persona designs to tailor different contexts, aligning with prior research that showed audio valence can evoke distinct emotions in listeners (Zentner et al., 2008; Juslin and Västfjäll, 2008; Tajadura and Västfjäll, 2008). Therefore, Study #2 also examined how the emotional component of the sounds used in audio personas influences social perception. Specifically, participants heard different types of sounds when interacting with each of the two research assistants. In the audio persona group, one research assistant had a positive audio persona, and the other had a negative audio persona.
@@
Nested factor: audio content. As Study #1 highlighted the wide range of potential audio content that could be used in audio persona designs, Study #2 considered the generalizability of the results by incorporating diverse sounds. Specifically, we included two different sounds for each valence condition (positive and negative), resulting in a collection of two positive and two negative sounds. For each trial, one positive and one negative sound were used.
@@ 5.1.3. Counterbalancing
We used a Latin Square design to counterbalance the audio content participants heard when each research assistant was present. This resulted in 4 audio options for one research assistant and 2 options for the other one (covering two valence levels). Furthermore, we counterbalanced the side of the room each research assistant entered, which determined the participant’s standing location as described in Sec. 5.1.2. Therefore, the sample size needed to be a multiple of (audio option for research assistant #1) (audio option for research assistant #2) (side of room entering) (condition groups) = to account for all combinations across both groups. Additionally, we randomized the order in which each research assistant entered and the side of the room where each speaker was located.
@@ 5.2.2. Task design
Participants and two research assistants took part in an activity called ”Something’s Different” (Krueger, 2009). We adapted this common icebreaker task to foster spontaneous interactions. In the activity, participants needed to identify a change (e.g., clothing or accessories) in each research assistant between two sessions. In each session, research assistants entered the room individually, walking back and forth in front of the participants. The changes to each research assistant were always on their name tags to remain consistency across participants: one switched from a yellow name tag to blue, and the other from a blue name tag to yellow.
5.2.3. Research assistants training
All participants interacted with the same two male, Asian research assistants, who wore identical clothing across study sessions to minimize the influence of attire on social perception. We chose to prescript the research assistants’ behavior in the activity to minimize confounding factors and isolate the specific impact of the audio personas on social perception. Both research assistants were trained to maintain neutral body language and facial expressions in the study, similar to (Tajadura-Jiménez et al., 2011). The research assistants were blinded about the conditions or sounds assigned in all sessions to ensure consistent behavior.
@@ 5.2.4. Familiarization phase
Before the icebreaker, participants walked around the study space to familiarize themselves. In the control group, participants heard sounds as they approached the speakers on each side of the room, as they would later during the task. In the audio persona group, they did not hear anything since no research assistants were present. Participants identified the presence of sounds and their sources before proceeding with the task.
@@ 5.3. Apparatus
Both body-anchored and object-anchored sounds were configured with a linear roll-off, with a minimum distance of 0.5 meters and a maximum distance of 2 meters. A 2-meter trigger distance was chosen based on audio persona designs from Study #1. Specifically, the audio persona group began hearing sound from a distance of 2 meters, with the volume increasing as the research assistants moved closer to them. The control group participants consistently heard the audio anchored to objects, regardless of the location of the research assistants. Details about sound characteristics are provided in the supplemental materials. Participants wore Bose QuietComfort Ultra Headphones in aware mode, allowing them to hear ambient sounds from the environment.
@@ 5.4. Procedure
Upon arriving at the lab, participants completed a consent form. Then, the experimenter introduced both research assistants to the participants by name and calibrated everyone’s head trackers based on their height. The participants were then introduced to the icebreaker task ”Something’s Different”, where their goal was to identify the change in each of the research assistants between two observation sessions. Before starting the task, researchers asked participants to walk around the study space to become familiar with it. Additionally, researchers showed participants the designated standing locations for observing the research assistants, as described in Sec. 5.1.
@@
After writing their answers, participants were instructed to walk to each research assistant and show them the results. The research assistant responded with a nod after reading. Once participants had shown their answers to both research assistants, they were asked to place their answer sheets back on the table and return to sit on one of the benches. The experimenter then entered to reveal the correct answer to the change and guided the participants to another room to complete a post-task questionnaire.
The study was approved by the Institutional Review Board (IRB # 71992). The full study took approximately 30 minutes, and all participants were compensated $15 for their time.

### v2

@@ Abstract.
We introduce Audio Personas, enabling users to ”decorate” themselves with body-anchored sounds in audio augmented reality. Like outfits, makeup, and fragrances, audio personas offer an alternative yet dynamic channel to augment face-to-face interactions. For instance, one can set their audio persona as rain sounds to reflect a bad mood, bee sounds to establish personal boundaries, or a playful ”woosh” sound to mimic passing by someone like a breeze. To instantiate the concept, we implemented a headphone-based prototype with multi-user tracking and audio streaming. Our preregistered in-lab study with 64 participants showed that audio personas influenced how participants formed impressions. Individuals with positive audio personas were rated as more socially attractive, more likable, and less threatening than those with negative audio personas. Our study with audio designers revealed that audio personas were preferred in public and semi-public-private spaces for managing social impressions (e.g., personality) and signaling current states (e.g., emotions).111Manuscript accepted to ACM Transactions on Computer-Human Interaction (TOCHI). This is the authors’ preprint version.
@@ 1. Introduction
While people can easily modify their social images online, face-to-face interactions lack such flexibility. Physical adornments like makeup and clothing remain relatively static, limited in their ability to convey changing social messages. As such, there is a growing trend toward integrating digital elements, traditionally limited to virtual interactions, into real-world settings. One prominent example is the use of visual filters, popularized through social media, are now adopted in mobile augmented reality (mobile AR). By anchoring visual effects such as face masks or virtual jewelries to one’s body, it alters how one is socially perceived by others (Häkkilä et al., 2017; Roinesalo et al., 2016).
@@
We introduce the concept of audio personas, body-anchored audio cues to augment social perception in face-to-face interactions. Here, the term persona refers to the aspect of someone’s character that is presented to others (Jung, 2014; Goffman et al., 2002), which can vary depending on the social context (Jung, 2014; Goffman et al., 2002; Leary and Kowalski, 1990; Tajfel, 1974). Sound effects are commonly used in online games to provide immediate feedback for user actions like hitting a barrier, or collecting a reward, enhancing player engagement (Liljedahl, 2011; Lipscombc, 2013; Kenwright, 2020). However, using audio in face-to-face social contexts is largely unexplored. In fact, psychological studies showed that exposure to emotion-evoking sounds can affect how people perceive others’ social attractiveness and emotions (May and Hamilton, 1980; Hanser et al., 2015; Logeswaran and Bhattacharya, 2009). For instance, women rated photos of men as more attractive when listening to positive music compared to when they heard negative music (May and Hamilton, 1980). With the rise of spatial audio, the ability to spatially anchor sounds to individuals—similar to how visual filters personalize appearance—creates a novel means of personalized expression through sounds.
@@
Figure 1 visualizes a few examples of audio personas. For instance, one can choose buzzing bee sounds to indicate personal boundaries, or a ”woosh” sound to mimic passing by like a breeze for fun and playfulness. One could also customize their footsteps to sound like bouncing balls or add a third footstep to express their love for gaming. Alternatively, keyboard-clicking sounds could be used to signal that the user is busy. To instantiate this concept, we first built a headphone-based prototype that utilizes multi-person tracking and multi-channel audio streaming. The user anchors sounds to their body and projects them to others in audio augmented reality.
@@
We conducted two user studies to explore the concept of audio personas. Study #1 (n = 64) examined whether audio personas are effective in influencing social perception, using a controlled, in-lab setup. Audio personas are designed to be body-anchored; therefore, as a between-subjects factor, we varied the spatial anchoring of sounds to assess the impact of this design choice. Participants interacted with two research assistants while hearing sounds either anchored to the assistants (i.e., audio personas) or to objects in the environment. As a within-subjects factor, we varied the valence of sounds to investigate how the emotional tone of audio personas influenced social perception. Overall, we found that participants were more likely to incorporate audio personas, body-anchored audio cues, into their impressions of others compared to the object-anchored audio. Moreover, participants perceived the person with a positive audio persona as more socially attractive, likable and less threatening than the person with a negative audio persona.
As Study #1 demonstrated that audio personas can influence social perception, Study #2 aimed to gain an understanding of use cases and design possibilities for audio personas. Through a design study with 8 audio designers, we identified intended use cases of audio personas in public and semi-public-private spaces as a way to manage social impressions (e.g., personality) and signal current states (e.g., emotions). We also elicited the creations of 48 audio personas from the designer participants and identified diverse design patterns in localization, triggers, dynamics and content of audio personas.
@@
Through a design study with audio designers (n = 8), we explored the social intentions and contexts for audio persona usage. This study also uncovered design possibilities for audio personas and identified key design patterns.
We discussed insights from the user studies and laid out design considerations and directions for future explorations of audio personas.
@@ 2.3. Dynamic Auditory Cues to Influence Social Perception
Our bodies generate sounds both within themselves (e.g., breathing, heartbeat) and through interactions with the environment (e.g., tapping sounds). Tajadura-Jiménez et al. introduced the concept of ”sonic self-avatar,” emphasizing that self-representations in mixed reality environments should incorporate body-centered sounds (Tajadura-Jiménez et al., 2018; Tajadura and Västfjäll, 2008). When body actions are sonified, the sonification can influence users’ mental representations of their bodies (Tajadura-Jiménez et al., 2015b; D’Adamo et al., 2024; Tonetto et al., 2014; Frid et al., 2016), perception of objects (Furfaro et al., 2015), and even gaits and motion (Tajadura-Jiménez et al., 2015a). For example, amplifying the high frequencies of footstep sound can create the self-perception of having a thinner body (Tajadura-Jiménez et al., 2015a) and sonifying body movement with descending pitch increases movement acceleration (Singh et al., 2024). These audio-based sensory feedback devices were further explored in the context of physical activity and rehabilitation (Ley-Flores et al., 2024; Turmo Vidal et al., 2024; Gomez-Andres et al., 2020).
@@ 3.2. Audio Rendering and Multi-channel Audio Streaming
4. Study #1: Can Audio Personas Affect Social Perception?
To explore whether audio personas are effective in influencing social perception, we conducted a preregistered study to examine key design components of audio personas—specifically audio anchoring and audio valence—in a controlled setting. 64 participants were invited to the lab to take part in an icebreaker activity with two research assistants.
@@ 4.1. Condition Design
Audio personas are proposed to be body-anchored. Thus, to understand the effect of audio personas on social perception, it’s essential to understand how this spatial anchoring plays a role.
Participants were divided into two groups. Half of the participants were assigned to the audio persona group (Figure 3a, 3b), in which participants heard sounds anchored to each research assistant. As the research assistant walked past participants (e.g., moving from left to right or vice versa), their audio persona shifted in locations corresponding to their movements.
In the control group, participants interacted with research assistants without hearing their audio personas. Instead, participants heard sounds anchored to objects (i.e., speakers) in the space. We chose to include sounds in the control group as previous research showed that the mere presence of audio can already influence social emotional perception and behavior (Tajadura-Jiménez et al., 2011; Logeswaran and Bhattacharya, 2009; Carroll and Young, 2005). This control design thus better examines the exact effect of ”body-anchored” audio. Anchoring sound to objects is also commonly used in audio augmented reality (McGookin et al., 2009; Yang et al., 2019; Sikora et al., 2018) and the auditory experience in this group is similar to hearing sounds played from a physical speaker. The research assistants’ movements would not affect the sound.
@@
Prior research has shown that audio valence can evoke distinct emotions in listeners (Zentner et al., 2008; Juslin and Västfjäll, 2008; Tajadura and Västfjäll, 2008). Therefore, Study #1 also examined how the emotional component of the sounds used in audio personas influences social perception. Specifically, participants heard different types of sounds when interacting with each of the two research assistants. In the audio persona group, one research assistant had a positive audio persona, and the other had a negative audio persona.
@@
Nested factor: audio content. To ensure generalizability of the results, we incorporated diverse sounds. Specifically, we included two different sounds for each valence condition (positive and negative), resulting in a collection of two positive and two negative sounds. For each trial, one positive and one negative sound were used.
@@ 4.1.3. Counterbalancing
We counterbalanced the audio valence (positive vs. negative) and the audio content (two options per valence) that participants heard during interactions with each research assistant. Additionally, we counterbalanced the side of the room from which each research assistant entered, which determined the participant’s standing position (as described in Section 5.1.2). Therefore, the sample size needed to be a multiple of (audio valence) × (audio content per valence) × (entry sides) × (research assistants) × (audio anchoring) = 32, to account for all combinations across both groups. Additionally, we randomized the order in which each research assistant entered and the side of the room where each speaker was located.
@@ 4.2.2. Task design
Participants and two research assistants took part in an activity called ”Something’s Different” (Krueger, 2009). We adapted this common icebreaker task to foster spontaneous interactions. In the activity, participants needed to identify a change (e.g., clothing or accessories) in each research assistant between two sessions. In each session, research assistants entered the room individually, walking back and forth in front of the participants. The changes to each research assistant were always on their name tags to remain consistent across participants: one switched from a yellow name tag to blue, and the other from a blue name tag to yellow.
@@ 4.2.3. Familiarization phase
Prior to the icebreaker, all participants completed a familiarization phase. This was intended to get all participants comfortable wearing the system, and also to prevent the control group from mistakenly attributing object-anchored sounds to the person. During this phase, participants were asked to walk around the space and identify whether any sounds were present and where they originated. The sounds were configured exactly as they would be during the icebreaker, and the control group heard sounds as they approached the corresponding objects. The research assistant was not present during this phase to ensure that participants encountered the research assistant for the first time during the main session. As a result, the audio persona group did not hear any sounds during this phase. All participants were able to identify whether sounds were present and, if so, determine their sources before proceeding with the icebreaker.
4.2.4. Research assistants
All participants interacted with the same two male, Asian research assistants, who wore identical clothing across study sessions to minimize the influence of attire on social perception. We chose to prescript the research assistants’ behavior in the activity to minimize confounding factors and isolate the specific impact of the audio personas on social perception. Both research assistants were trained to maintain neutral body language and facial expressions in the study, similar to (Tajadura-Jiménez et al., 2011). They were instructed to walk at a steady pace, avoid expressive gestures and facial expressions, and maintain forward-facing eye gaze. An experimenter provided real-time feedback during repeated practice to ensure consistent and neutral behavior. In the study sessions, research assistants were blinded about the conditions assigned to the participants.
@@ 4.3. Apparatus
Both body-anchored and object-anchored sounds were configured with a linear roll-off, with a minimum distance of 0.5 meters and a maximum distance of 2 meters. A 2-meter trigger distance was chosen following E.T. Hall’s ”social distance” range (Hall, 1963). Specifically, the audio persona group began hearing sound from a distance of 2 meters, with the volume increasing as the research assistants moved closer to them. The control group participants consistently heard the audio anchored to objects, regardless of the location of the research assistants. Details about sound characteristics are provided in the supplemental materials. Participants wore Bose QuietComfort Ultra Headphones in aware mode, allowing them to hear ambient sounds from the environment.
@@ 4.4. Procedure
Upon arriving at the lab, participants completed a consent form. Then, the experimenter introduced both research assistants to the participants by name and calibrated everyone’s head trackers based on their height. The participants were then introduced to the icebreaker task ”Something’s Different”, where their goal was to identify the change in each of the research assistants between two observation sessions. Before starting the task, researchers asked participants to walk around the study space to become familiar with it. Additionally, researchers showed participants the designated standing locations for observing the research assistants, as described previously.
@@
After writing their answers, participants were instructed to walk to each research assistant and show them the results. The research assistant responded with a nod after reading. Once participants had shown their answers to both research assistants, they were asked to place their answer sheets back on the table and return to sit on one of the benches. During this phase, behavioral measures including participants’ physical distance from each research assistant and their seat selection were recorded. Participants could still hear the sounds anchored to the person or object if they were within the trigger distance. At the end, experimenters entered to reveal the correct answer to the change and guided the participants to another room to complete a post-task questionnaire.
The study was approved by Stanford Institutional Review Board (IRB # 71992). The full study took approximately 30 minutes, and all participants were compensated $15 for their time.